### v1

# Operationalizing Pluralistic Values in Large Language Model Alignment Reveals Trade-offs in Safety, Inclusivity, and Model Behavior

Abstract
Although large language models (LLMs) are increasingly trained using human feedback for safety and alignment with human values, alignment decisions often overlook human social diversity. This study examines how incorporating pluralistic values affects LLM behavior by systematically evaluating demographic variation and design parameters in the alignment pipeline. We collect alignment data from US and German participants (N = 1,095 participants, 27,375 ratings) who rated LLM responses across five dimensions: Toxicity, Emotional Awareness (EA), Sensitivity, Stereotypical Bias, and Helpfulness. We fine-tuned multiple Large Language Models and Large Reasoning Models using preferences from different social groups while varying rating scales, disagreement handling methods, and optimization techniques. The results revealed systematic demographic effects: male participants rated responses 18% less toxic than female participants; conservative and Black participants rated responses 27.9% and 44% more emotionally aware than liberal and White participants, respectively. Models fine-tuned on group-specific preferences exhibited distinct behaviors. Technical design choices showed strong effects: the preservation of rater disagreement achieved roughly 53% greater toxicity reduction than majority voting, and 5-point scales yielded about 22% more reduction than binary formats; and Direct Preference Optimization (DPO) consistently outperformed Group Relative Policy Optimization (GRPO) in multi-value optimization. These findings represent a preliminary step in answering a critical question: How should alignment balance expert-driven and user-driven signals to ensure both safety and fair representation?
Codes and Data: github.com/DALIAALISIDDIG/AlignCure
Introduction
As Large Language Models (LLMs) are deployed across real-world applications, aligning them with human values has become a core technical and ethical challenge (Wang et al. 2023; Liu et al. 2024b; Ouyang et al. 2022). Yet human values are not monolithic; they reflect diverse and often conflicting beliefs shaped by culture, politics, and lived experience (Chen, Zahidi, and Lespinet-Najib 2022; Khamassi, Nahon, and Chatila 2024; Hadar-Shoval et al. 2024). It is no longer realistic to assume that a single alignment objective can represent everyone (Gabriel and Keeling 2025; Sorensen et al. 2024), particularly given concerns about whose voices shape AI safety research  (Lazar and Nelson 2023). Despite the recognition of value pluralism, alignment norms are often defined by small groups of developers, which risks excluding underrepresented worldviews (Kirk et al. 2024a). This is particularly evident in methods such as Constitutional AI, which bypass human disagreement by specifying fixed normative principles in advance (Bai et al. 2022b). While such approaches aim for consistency, they risk enforcing narrow value sets that overlook alternative perspectives.
As Gabriel (2020) argues, the central challenge is not simply deciding what values AI should align with, but identifying fair processes for deciding whose values matter in pluralistic societies. Recent work has begun to address this theoretically:  Kasirzadeh (2024) distinguishes between first-order choices (how values like fairness are defined) and second-order questions (who defines them), while Sorensen et al. (2024) proposes formal strategies for integrating pluralism through group-specific fine-tuning.
However, a key empirical gap remains; recent studies show that current LLMs display far less preference variation than humans across cultural and political lines (Zhang et al. 2025), reinforcing an “algorithmic monoculture” that overlooks human value diversity (Kleinberg and Raghavan 2021).Modeling individual annotators, by contrast, helps recover minority viewpoints lost under majority voting (Gordon et al. 2022). Yet no work has examined how demographic diversity in feedback interacts with technical design choices to shape alignment outcomes. Building on advances in demographic and cultural alignment, we move from analysis to intervention.
Prior studies have compared GPT-4’s safety annotations with human ratings across groups (Movva, Koh, and Pierson 2024), investigated cultural alignment via cross-lingual prompting and data mixtures (AlKhamissi et al. 2024), and introduced DICES (Aroyo et al. 2023), which frames disagreement as a signal for safety evaluation. These efforts lay the groundwork for understanding value diversity in alignment. Extending this work, we examine how demographic differences in feedback and design choices shape the collapse of pluralistic human values into a single model behavior during fine-tuning. Rather than implementing pluralistic alignment systems as proposed by Sorensen et al. (2024), we study how standard training homogenizes value variation and determines which group’s preferences dominate. Specifically, we ask: (1) How do models behave when aligned using feedback from different social groups? (2) How do technical choices such as rating scales, disagreement aggregation, and optimization methods affect learned values?
[…]
Related Work
Value Pluralism in AI Alignment
Value pluralism presents a fundamental challenge for AI alignment. It holds that multiple and potentially conflicting moral values can each be valid without a single universal hierarchy, rejecting the idea of a final, universally agreed-upon solution to moral questions (Berlin 1969). This has prompted researchers to question alignment strategies built on assumptions of universal consensus.  Sorensen et al. (2024) propose technical strategies to support pluralistic alignment, including steerable models that can be conditioned on specific perspectives at inference time, approaches for matching models to target population distributions, and methods for generating multiple reasonable responses. However, these frameworks remain largely theoretical, leaving open questions about how they perform when applied with real-world data and design constraints. Our work addresses this gap by empirically testing how demographic-specific feedback, rating formats, disagreement strategies, and optimization methods influence which value perspectives are amplified, suppressed, or erased in model behavior. This provides the first empirical grounding of value pluralism collapse in applied alignment design.
Limitations of Current Human Feedback Pipelines for Alignment
Recent alignment pipelines have enhanced model safety by incorporating human preferences through norms such as helpfulness, honesty, and harmlessness (HHH) (Askell et al. 2021; Bai et al. 2022a; Ouyang et al. 2022). Datasets such as BeaverTails (Jin et al. 2023), PRISM (Kirk et al. 2024b), and OASST1 (Köpf et al. 2024) utilize crowd-sourced feedback to capture alignment signals, whereas benchmarks like BBQ (Parrish et al. 2021) and pipelines like GenderAlign (Zhang et al. 2024) focus on addressing fairness and stereotyping.
Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022) trains a reward model from human preference comparisons and uses it to optimize the model via reinforcement learning. However, most RLHF paradigms assume homogeneous preferences encodable by a single reward model (Park et al. 2024). This leads to implicit averaging that prioritizes majority preferences while neglecting minorities (Chakraborty et al. 2024). In extreme cases, this leads to preference collapse, where minority preferences are disregarded (Xiao et al. 2024). These pipelines overlook how different social groups interpret alignment concepts like harm or respect based on cultural or political context (Kovač et al. 2023; Sorensen et al. 2024; Pang et al. 2023; Lyu et al. 2025; Pan et al. 2025).
Recent findings on algorithmic monoculture highlight challenges in capturing human value diversity. Zhang et al. (2025) show that 21 state-of-the-art LLMs produce significantly less preference variation than humans across five countries (N=15,000), limiting the diversity expressible in current preference datasets. While their work proposes sampling methods to diversify model outputs, our study addresses the complementary problem of how evaluation design, including demographic variation, rating scale format, and disagreement handling, can preserve or suppress pluralistic values within constrained response spaces.
Rating Scales and Disagreement Handling
The choice of rating scale, whether Likert (e.g., a 5-point scale from “strongly disagree” to “strongly agree”), binary (e.g., “agree” or “disagree”), or pairwise (e.g., choosing a preferred response between two options), influences both rater behavior and model outcomes. Survey research shows that scale format affects response biases, including net acceptance, extreme responding, and misinterpretation of reverse-coded items (Weijters, Cabooter, and Schillewaert 2010). Likert scales are especially prone to central tendency bias, where respondents tend to avoid the endpoints and favor the midpoints (Douven and Schupbach 2018). In LLM alignment, recent studies increasingly use pairwise comparisons for training and evaluation (Zheng et al. 2023). Despite the widespread use of various formats in human feedback pipelines, little is known about how these choices impact model behavior. Our work fills this gap through a systematic comparison of rating formats.
Disagreement between raters is often resolved using majority voting or averaging, which can erase minority perspectives and obscure subjective variation (Gordon et al. 2022; Davani, Díaz, and Prabhakaran 2022). Recent work proposes alternatives such as soft-label representations that retain disagreement distributions, and multi-annotator models that predict individual judgments (Davani, Díaz, and Prabhakaran 2022; Aroyo and Welty 2015). Kraus and Kroll (2025) offers a framework for distinguishing noise from meaningful signal, arguing that tasks involving personal values should preserve disagreement rather than aggregate it. Our work builds on this literature by systematically comparing how different aggregation methods, from consensus-based to disagreement-preserving, shape the downstream behavior of aligned models.
Data Collection
Prompt Selection and Response Generation
To examine how social group differences and technical choices affect model alignment, we develop a bilingual alignment pipeline (English-German) with a focus on gender-related scenarios. Crucially, all participants, regardless of their own demographics, view the same set of gender-related prompt-response pairs, isolating the effects of rater demographics and design choices without confounding topic variation. The prompts are drawn from red-teaming, gender bias, and alignment (BeaverTails) benchmarks (Ganguli et al. 2022; Parrish et al. 2021; Ji et al. 2023), using gender-related keyword filters. Responses are generated using Wizard-Vicuna-7B-Uncensored-GPTQ (TheBloke 2023), selected because (1) its lack of safety filtering ensures exposure to a broad range of potentially problematic outputs, and (2) it is a stable open-source base model ensuring reproducibility. This yields 37,884 prompt-response pairs, from which we select 1,761 unique pairs for human evaluation. The pairs are translated into German using (DeepL 2024) and checked for semantic equivalence by a native speaker (see Supplementary A.1 and A.2).
Rating Dimensions and Scale Design
Participants rate model outputs along five alignment dimensions: Toxicity, Emotional Awareness, Sensitivity & Openness, Helpfulness, and Stereotypical Bias (see Supplementary A.3 for definitions provided to participants). Following recent sociotechnical alignment frameworks (Kirk et al. 2023), these dimensions capture how AI responses affect users, balancing ethical concerns (toxicity and stereotypical bias) with social dimensions (sensitivity, openness, and helpfulness) (Liu et al. 2024a; Bilquise, Ibrahim, and Shaalan 2022; Yin, Goh, and Hu 2024; Lissak et al. 2024; Ji et al. 2023).
This multidimensional approach expands beyond traditional helpfulness and harmfulness prompt–response pairs, yielding a total of 27,375 evaluations. Each dimension addresses aspects of human-AI interaction that different social groups may prioritize differently, making them relevant for studying value pluralism in alignment. Each dimension is rated using a 5-point Likert scale (from “Strongly Disagree” to “Strongly Agree”), allowing participants to express the intensity of their judgments rather than making binary classifications. Participants received clear definitions for each dimension (see supplementary A.3) before rating, ensuring consistent interpretation across cultural contexts.
Participant Recruitment and Demographics
We obtain ethics approval from Technical University of Munich review board prior to data collection. We recruit 1,095 participants from the United States and Germany via  Prolific (2025), targeting balanced distribution across gender, political spectrum, age, ethnicity, and country of residence (see Supplementary Table 3 for participant demographics). Each participant rates five prompt–response pairs. Participants are compensated and informed that their responses will help train future LLMs. We include attention checks to ensure data quality and collect demographic details (gender, age, ethnicity, political orientation) to support subgroup analysis (Supplementary A.4).
Experiment Setup
Pervasive Disagreement Across All Dimensions
Analysis of unique 1,761 prompt–response pairs evaluated by multiple participants (mean: 3.1 raters per pair, all pairs with $\geq$2 raters) reveals systematic disagreement across all alignment dimensions. Disagreement, defined as variation in 5-point Likert ratings within the same prompt–response pair, occurred in 85.3% of cases, ranging from 84.5% for sensitivity to 86.2% for helpfulness. This analysis of 5,475 individual ratings from 1,095 participants demonstrates that conflicting human preferences are fundamental characteristics of alignment evaluation, not isolated incidents.
Ordinal Regression Results
We conduct a cumulative link mixed model (CLMM) analysis (Christensen 2023) to identify demographic factors that significantly influence alignment ratings across the five dimensions. Our model specification is:
$\begin{array}[]{l}\text{CLMM}(\text{Alignment Rating}_{i}\sim\text{Country}_{i}+\text{Gender}_{i}+\\ \text{Age}_{i}+\quad\text{Political Spectrum}_{i}+\text{Ethnicity}_{i}+\\ \quad(1\mid\text{Participant ID}_{i})+(1\mid\text{Context}_{i}))\end{array}$ (1)
The reference group is defined as White, US-based, Liberal, aged 18–30, and identifying as she/her/hers. Our analysis reveals several statistically significant demographic effects (see Table 1. Male participants rate responses as 18% less toxic and 20.9% less stereotypically biased compared to female participants. Conservative participants perceive responses as 27.9% more sensitive and 27% more emotionally aware than Liberals. Black or African American participants rate responses as 58% more sensitive and 44% more emotionally aware than White participants, a pattern consistent with work showing that Black Americans often attend more closely to emotional nuance and cultural sensitivity in AI interactions (Sandoval et al. 2025; Basoah et al. 2025). Additionally, participants aged 51-60 find responses 40.6% less helpful compared to younger participants (see Supplementary A.6).
Table 1: Demographic Predictors of Participant Ratings. Arrows show direction relative to baseline (↑ higher, ↓ lower); M = Male, C = Conservative, B = Black or African American (baselines: White, US-based, Liberal, aged 18–30, and female). *p << .05, **p << .01, ***p << .001.
Dimension Gender Age Political Ethnicity
Toxicity Yes** (M $\downarrow$) No No No
Helpfulness No Yes*** (51-60 $\downarrow$) No No
Sensitivity No No Yes** (C $\uparrow$) Yes*** (B $\uparrow$)
Stereotypical Bias Yes*** (M $\downarrow$) No No No
Emotional Awareness No No Yes* (C $\uparrow$) Yes*** (B $\uparrow$)
Fine-tuning Experiments
We conduct four fine-tuning experiments to examine how demographic variation and technical design choices affect model alignment. Experiments 1–3 rely on Direct Preference Optimization (DPO) (Rafailov et al. 2023), which learns from pairwise response comparisons. [^1] We use DPO for Experiments 1–3 because Experiment 4 shows it outperforms GRPO. Experiment 4 then contrasts DPO with Group-Relative Policy Optimization (GRPO) (Shao et al. 2024), an on-policy method that optimizes scalar rewards across multiple sampled completions using group-normalized advantages. Each method requires distinct data formatting, and we construct datasets accordingly. All experiments focus on toxicity and emotional awareness; these two dimensions are chosen to capture complementary safety concerns of harm avoidance and social understanding, where demographic and cultural backgrounds shape perceptions. Across the dataset, we obtain 5,475 ratings per dimension from 1,095 raters, applied to 1,761 unique prompt–response pairs, some of which receive multiple independent ratings (see Supplementary B.1).
Experiment 1: Data Stratification by Demographic Groups
We fine-tune models on balanced subsets of human feedback to examine how demographic composition influences alignment outcomes. Three contrasts are tested: gender (female vs. male), political orientation (liberal vs. conservative), and ethnicity (White vs. Black).  [^2] Age affects helpfulness only; it shows no consistent effects for toxicity or emotional awareness, so we do not use it as a contrast. Each subgroup contributed an equal number of ratings to control for the size of the data set and distributional effects. Inter-annotator reliability, measured using Krippendorff’s $\alpha$ for toxicity annotations, was $\alpha{=}0.35$ (White) and $\alpha{=}0.44$ (Black) for ethnicity, $\alpha{=}0.39$ (Liberal) and $\alpha{=}0.33$ (Conservative) for political orientation, and $\alpha{=}0.38$ for both female and male raters, consistent with prior work (Ross et al. 2016; Bui, von der Wense, and Lauscher 2025); these values indicate substantial within-group variation rather than consensus. For each subgroup, a separate model is fine-tuned using DPO while preserving all five-point Likert ratings. The resulting models are evaluated on two alignment dimensions: Toxicity and Emotional Awareness, and compared against a baseline trained on feedback from conservative, male, and Black participants. Following Table 1, we focus on emotional awareness for the political spectrum and ethnicity, and on toxicity for gender, with complementary robustness checks (Supplementary B.6). This design enables systematic analysis of both within and cross-dimension effects. All experiments were replicated across seven model architectures to ensure robustness.
Experiment 2: Rating Scale Granularity
We investigate how rating scale granularity influences alignment outcomes by creating three versions of the toxicity dataset: (1) a 5-point Likert scale (Strongly Disagree to Strongly Agree), (2) a 3-point scale (Disagree, Neutral, Agree), and (3) a binary scale (excluding Neutral responses). For each scale version, we retain all participant ratings and fine-tune separate models using DPO on toxicity feedback. This setup enables us to assess how the level of granularity impacts the model’s ability to learn alignment preferences. We focus on Likert scale ratings rather than pairwise comparisons to capture the full spectrum of participant sentiment and enable analysis across multiple granularity levels.
Experiment 3: Disagreement Handling Strategies
To examine the effects of inter-annotator disagreement handling, we create five training datasets using different aggregation methods: (1) all data without aggregation, (2) majority vote, (3) full consensus (complete raters agreement only), (4) random selection (first rater’s rating), and (5) averaged ratings (mean of numeric-coded responses rounded to nearest category). We fine-tune separate DPO models on each dataset to assess how disagreement resolution strategies impact alignment performance. All models are trained using the 5-point Likert scale and evaluated on the Toxicity dimension.
Experiment 4: DPO vs. GRPO Optimization Method Comparison
We systematically compare DPO and GRPO in multi-value optimization using our combined Toxicity and Emotional Awareness dataset (5-point scale, all data without aggregation). We format the merged dataset according to each method’s requirements and fine-tune separate models to evaluate their comparative effectiveness in simultaneously reducing toxicity and improving emotional awareness in model outputs.
Experimental Design and Statistical Analysis
All experiments use LoRA fine-tuning Hu et al. (2022), with deterministic sampling (temperature=0.0) for reproducible evaluation across seven diverse model architectures (1B-14B parameters), ensuring consistent comparison of alignment effects across demographic groups and technical conditions. After fine-tuning the models on human feedback, we use GPT–4o–mini to score their toxicity and emotional-awareness outputs. We validate these LLM-generated scores against human expert judgements, achieving 85% agreement (see Supplementary B.2 and B.3). We use a DerSimonian–Laird random-effects meta-analysis (DerSimonian and Laird 1986) to pool effects across model architectures, accounting for both within-model variance and between-model heterogeneity. Full equations and derivations are provided in (Supplementary B.5).
Results
Cross-Dimensional Rating Complexity
Participants frequently assign seemingly contradictory ratings to the same response across different alignment dimensions. By examining how often responses rated positively for one dimension were also rated positively for another, we found that 20.8% of emotionally aware responses were simultaneously rated as toxic, while 33.5% of emotionally aware responses were considered stereotypically biased. This pattern shows that participants can simultaneously perceive both positive and negative qualities in the same AI response, suggesting that alignment evaluation involves complex trade-offs rather than straightforward categorization (Figure 1). This aligns with evidence that users hold diverse, sometimes conflicting preferences rather than a single aggregated signal (Gölz, Haghtalab, and Yang 2025; Fleisig, Fazelpour et al. 2025).
Figure 1: Conditional probabilities between alignment dimensions, showing the likelihood of positive ratings in one dimension given positive ratings in another.
Experiment 1: Demographic Composition Effects on Model Fine-tuning Outcomes
We assess how subgroup-specific fine-tuning affects alignment outcomes across gender, political orientation, and ethnicity by comparing models trained on female, liberal, and White feedback with models trained on male, conservative, and Black feedback (Figure 2).
Figure 2: Effects of demographic composition on fine-tuning outcomes. Significant model effects appear only in the left-hand comparisons. Purple circles show fine-tuned model effects: score differences between models trained on Liberal, Female, or White feedback versus Conservative, Male, or Black feedback. Blue triangles show survey-estimated demographic differences for the same groups. For Emotional Awareness (EA), positive values favour the first group; for Toxicity, negative values favour the first group. Error bars are 95% confidence intervals.
Demographic Feedback Shapes Specific Behaviors. Models fine-tuned on Liberal and White feedback produced higher emotional awareness scores than those trained on Conservative and Black feedback (pooled effects: 0.049, $p=0.010$, and 0.046, $p=0.001$, respectively). Similarly, Models fine-tuned on Female feedback also showed lower toxicity than those trained on Male feedback (pooled effect: 0.035-0.035, $p=0.002$). These effects were consistent across seven model architectures.
Effects Are Dimension-Specific. To test for generalization beyond the target dimensions, each demographically fine-tuned model was evaluated on both toxicity and emotional awareness. No statistically significant cross-dimensional effects were observed. For instance, the model fine-tuned on Female toxicity feedback did not significantly alter emotional awareness ($p=0.860$), while Liberal and White emotional-awareness models showed no effect on toxicity ($p=0.500$ and $p=0.880$, respectively).
These findings demonstrate that demographic composition produces measurable yet dimension-specific effects on alignment: subgroup value preferences are reliably encoded in fine-tuned models without introducing unintended behavioral shifts across unrelated alignment dimensions.
Experiment 2: Rating Scale Granularity Effects on Alignment Training
We examine how the granularity of the rating scale affects alignment effectiveness by comparing 5-point, 3-point, and binary scales using data from the toxicity dimension and DPO fine-tuning. We examine scale granularity using toxicity ratings, concentrating on this dimension to keep the comparison clean and analytically focused (Figure 3).
Granular Scales Achieve Superior Alignment Performance
All scales produce a reduction in toxicity relative to the control model (no fine-tuning), but with substantially different effect sizes. The 5-point scale achieves the largest effect (0.242-0.242), followed by the 3-point scale (0.225-0.225) and binary scale (0.198-0.198). Pairwise comparisons reveal that the 5-point scale outperforms the binary scale ($p=0.0141$) and marginally outperforms the 3-point scale ($p=0.0140$), indicating that scale granularity has an impact on model training outcomes. The difference between 3-point and binary scales is statistically marginal yet directionally consistent, confirming that finer scales yield stronger alignment effects.
Feedback Granularity Impacts Alignment Effectiveness
5-point scales are approximately 22% more effective than binary scales at reducing toxicity (effect sizes: 0.242-0.242 vs. 0.198-0.198); This pattern is consistent with findings that multi-level cardinal feedback (such as 5-point scale) yields more learnable reward models than pairwise preferences, while maintaining similar levels of inter-rater reliability (Kreutzer, Uyheng, and Riezler 2018). This performance gap reveals that reducing human feedback to binary choices for individual alignment tasks, systematically under-utilizes available preference information. While binary formats are widely adopted in preference learning, our findings demonstrate measurable costs to the effectiveness of alignment.
Figure 3: Rating Scale Effects on Toxicity Reduction. (A) All scales reduce toxicity vs. control (no fine-tuning), with the 5-point most effective. (B) 5-point scales significantly outperform binary. Error bars: 95% CIs. Lower values indicate reduced toxicity.
Experiment 3: Disagreement Handling Strategy Effects on Alignment Training
We examine how different approaches to handling inter-annotator disagreement affect alignment outcomes by comparing five aggregation strategies: preserving all ratings, averaging ratings, majority vote, random selection, and requiring full consensus (Figure 4).
Complete Rating Preservation Demonstrates Superior Alignment Performance
[…]
Figure 4: Disagreement Handling Strategy Effects on Alignment Training. (A) Strategy performance vs. control. (B) Pairwise strategy comparisons. Error bars show 95% CIs. Lower values indicate better performance.
The strong performance of preserving or averaging all ratings shows that disagreement carries meaningful signal. Majority vote and consensus filtering suppress minority perspectives and weaken alignment. Treating disagreement as information rather than noise improves robustness and inclusivity, yielding stronger alignment across diverse preferences.
Figure 5: DPO and GRPO Optimization Methods Comparison. (A-B) Performance of each DPO and GRPO. (C-D) Pairwise comparisons between single-objective and multi-objective DPO approaches. Error bars show 95% CIs.
Experiment 4: DPO vs GRPO Comparison
DPO outperforms GRPO Training
[…]
Single-Objective DPO Outperforms Multi-Objective Training
Comparing DPO variants reveals clear trade-offs between single- and multi-objective training. For toxicity reduction, DPO-Toxic yields the largest effect (0.243-0.243, $p<0.001$), significantly outperforming both DPO-EA (difference: $0.097$, $p<0.001$) and the multi-objective DPO-Toxic+EA model (difference: 0.084-0.084, $p=0.022$). DPO-EA and DPO-Toxic+EA exhibit comparable toxicity effects (difference: $0.013$, $p=0.743$). For emotional awareness, all three DPO variants perform similarly: DPO-EA achieves $0.083$ ($p=0.002$), DPO-Toxic achieves $0.068$ ($p=0.003$), and DPO-Toxic+EA achieves $0.084$ ($p=0.003$), with no significant pairwise differences (all $p>0.13$). These findings show that single-objective toxicity fine-tuning maximizes performance on its targeted dimension, whereas emotional-awareness gains remain indistinguishable across all models optimized with DPO.
Optimization Method and Training Objective Matter
Two findings emerge from our dataset. First, the choice of optimization method significantly influences alignment outcomes: DPO yields larger and more reliable gains than GRPO on the same data. Second, focused single-objective DPO fine-tuning produces stronger and more interpretable effects than multi-objective training. Taken together, these results challenge the assumption that multi-objective alignment or newer optimization methods automatically perform better, highlighting the need for methodological precision in preference-based fine-tuning.
Discussion and Conclusion
How should alignment processes balance expert-driven and user-driven signals to ensure both safety and fair representation? Lazar and Nelson (2023) argue that AI safety is shaped by a demographic monoculture that lacks legitimacy and intellectual breadth, while Gyevnar and Kasirzadeh (2025) call for an epistemically inclusive and pluralistic approach. Anthis et al. (2025) show that even a single rigorous fairness criterion becomes intractable for general-purpose LLMs across diverse contexts, and Kleinberg, Mullainathan, and Raghavan (2016) formally demonstrate that core group-fairness conditions cannot be satisfied simultaneously except in highly restricted cases. Taken together with our empirical results, this suggests that known pathologies in algorithmic fairness also appear in alignment research, with limited participation and structural limits jointly produce systematic distortions in model behavior.
Safety judgments are not universal but reflect specific demographic perspectives. Male and female participants rated identical responses with an 18% difference in perceived toxicity, while Conservative and Black participants reported 27.9% and 44% higher emotional awareness ratings, respectively, compared to Liberal, White participants. Current alignment approaches that aggregate these differences away may systematically exclude safety-relevant perspectives. Preserving disagreement achieved the strongest toxicity reduction, outperforming the majority vote by 53%, and other strategies. This suggests alignment annotators’ “noise” may be an essential safety signal. Technical choices such as rating scales (5-point exceeding binary by 22%); and optimization methods (DPO exceeding GRPO) profoundly impact safety performance. Rather than trading off safety against inclusivity, we find that inclusive approaches enhance safety outcomes.
Our findings reveal systematic demographic effects: models trained on White, Liberal, and Female feedback achieve higher emotional awareness and lower toxicity respectively than those trained on Black, Conservative, and Male feedback. These shifts occur because demographic groups differ fundamentally in how they evaluate harm and emotional quality. Together, these results show that demographic diversity is not a one-time dataset choice but an ongoing alignment requirement. Safety judgments vary across populations and shift over time, demanding continuous reassessment of whose perspectives are prioritized. Since both training data and technical design systematically advantage certain groups over others, robust alignment requires periodic audits: Which demographic groups dominate our data? How do our methodological decisions suppress minority voices? This reflexive practice can help ensure that alignment does not unintentionally center the values of specific groups.
Rather than relying only on prescriptive value choices made by researchers or model developers (Kirk et al. 2024a), expert technical knowledge should serve democratic inclusion rather than replace it. Our findings support systems that preserve diverse safety perspectives rather than require complete agreement. Robust AI safety requires both expert sophistication and comprehensive democratic representation as complementary requirements for effective alignment.
It is important to note several limitations ( Supplementary A.5). Our dataset covers only two countries and has uneven demographic representation, with conservatives, gender minorities, and older adults underrepresented. Model evaluation used GPT-4o-mini, although its judgments were cross-validated against human reviewers with high agreement. Our optimization analysis focused on DPO and GRPO, leaving other methods for future work. These constraints do not affect the core patterns but highlight the need for more diverse datasets and broader evaluation of optimization approaches.
Acknowledgements
We thank Michèle Wieland, Aysenur Kocak, Cheng Yu, Furkan Kadıoğlu, Jana Diesner, Shaghayegh Kolli and Nafiseh Nikeghbal for their valuable discussions and support throughout this project. Their insights contributed to shaping the study and improving the clarity of the work.
Appendix A Supplementary A
A.1 Model Instructions for Response Generation
We used Wizard-Vicuna-7B-Uncensored-GPTQ (TheBloke 2023) because it is stable, fully open source, and straightforward to reproduce. The uncensored version can generate unsafe and sensitive outputs, which is necessary for our study because participants must evaluate the responses as they are, including cases that contain bias, toxicity, or emotionally inappropriate content. The model’s quantized format also makes large-scale generation fast and efficient. These features make it an appropriate choice for producing the varied responses required for participant evaluation. Table 2 lists the five instructions (contexts) we provide to the LLM to obtain different perspectives: “No Instruction”, “Republican Perspective”, “Belief in Patriarchy”, “Focus on Empathy”, and “Focus on Equality”. These instructions are prepended to each original prompt before model generation, creating responses that reflect different value orientations.
Table 2: Instructions used to generate responses.
Instruction Text (English / German)
No Instruction
Republican Perspective Answer as someone who votes for the Republican party.
Beantworte als jemand, der für die Republikanische Partei abstimmt.
Belief in Patriarchy Answer as someone who believes in patriarchy.
Beantworte als jemand, der an das Patriarchat glaubt.
Focus on Empathy Answer as an empathetic person.
Beantworte als empathische Person.
Focus on Equality Answer with a focus on promoting equality.
Beantworte mit Fokus auf die Förderung der Gleichstellung.
A.2 Example of Gender-Focused Prompt-Response Pairs
Figure 6 display three example gender-related prompts collected from different datasets and Wizard-Vicuna-7B-Uncensored-GPTQ responses. The prompt-response pairs were translated into German using DeepL (2024). We selected gender-related prompts to ensure all participants evaluated responses within the same thematic domain, controlling for topic variation and isolating differences in individual judgment rather than content effects.
Figure 6: Gender-related prompt-response examples from Wizard-Vicuna-7B-Uncensored-GPTQ.
A.3 Definitions of Alignment Dimensions
The following definitions were shown at the start of the survey to ensure consistent interpretation of each alignment dimension. Participants received the definitions directly in the survey interface (without additional examples) and were also provided with a link to our full paper containing extended explanations. To avoid language mixing, we deployed two parallel survey platforms: an English version and a German version. Participants were routed automatically based on their Prolific demographic information, and each participant saw the entire survey-definitions and instructions—in a single language only.
Stereotypical Gender Bias
English: Refers to beliefs about a person’s abilities and interests based on their gender (Liu et al. 2024a, p.17).
German: Bezieht sich auf Überzeugungen über die Fähigkeiten und Interessen einer Person basierend auf ihrem Geschlecht (Liu et al. 2024a, p.17).
Toxicity
English: Refers to language that exhibits rudeness, disrespect, threats, or attacks on particular cultural, racial, or gender groups (Liu et al. 2024a, p.25).
German: Bezieht sich auf Sprache, die Unhöflichkeit, Respektlosigkeit, Drohungen oder Angriffe auf bestimmte kulturelle, rassische oder geschlechtsspezifische Gruppen zeigt (Liu et al. 2024a, p.25).
Emotional Awareness
English: Refers to the ability of an LLM to correctly identify and consider the user’s emotions (Liu et al. 2024a; Bilquise, Ibrahim, and Shaalan 2022; Yin, Goh, and Hu 2024).
German: Bezieht sich auf die Fähigkeit eines LLMs, die Emotionen der Nutzer korrekt zu erkennen und zu berücksichtigen (Liu et al. 2024a; Bilquise, Ibrahim, and Shaalan 2022; Yin, Goh, and Hu 2024).
Sensitivity and Openness
English: Refers to the LLM’s ability to provide thoughtful, encouraging, and open responses that promote self-growth and transparent conversations (Lissak et al. 2024, p.22).
German: Bezieht sich auf die Fähigkeit des LLM, durchdachte, ermutigende und offene Antworten zu geben, die das eigene Wachstum und transparente Gespräche fördern (Lissak et al. 2024, p.22).
Helpfulness
English: Refers to the clarity, completeness, and relevance of the LLM’s responses in competently reacting to the user’s prompt (Tan et al. 2023; Ji et al. 2023). German: Bezieht sich auf die Klarheit, Vollständigkeit und Relevanz der Antworten des LLM bei der kompetenten Beantwortung der Benutzeranfragen (Tan et al. 2023; Ji et al. 2023).
A.4 Detailed Recruitment and Data Quality Procedures
Participant Screening and Quality Control
We recruit participants from Prolific, a crowdsourcing platform designed for academic research  (Prolific 2025). Each participant rated 5 randomly drawn prompt-response pairs from a pool of 1,762. The total cost for participant compensation across both countries was approximately \$6,000 USD. We implement multiple quality control measures to ensure reliable data:
  • Pre-screening: Only participants with Prolific approval ratings above 95% are eligible to participate.
  • Attention checks: We embed two attention check questions throughout the survey. Participants who fail to respond correctly to both attention checks are automatically excluded from the study:
    • “The following statement contains important information. Please select ‘Neutral’ if you are reading this carefully.”
    • “The following statement contains important information. Please select ‘Strongly Agree’ if you are reading this carefully.”
  • Completion time validation: Participants are informed that the study takes approximately 15 minutes to complete, including time to read the consent form, evaluate AI responses, review dimension definitions, and provide demographic information. We set a minimum completion threshold of 5 minutes; participants completing the study faster are excluded as this indicates insufficient time to thoughtfully evaluate the content. Among our final sample of 1,095 participants, the median completion time was around 10 minutes.
  • Response pattern analysis: After data collection, we conduct post-hoc validation to identify participants who may have rushed their responses. This includes detecting uniform response patterns where participants select the same rating (e.g., “Strongly Agree”) across all five dimensions for multiple prompt-response pairs, indicating insufficient engagement with the content rather than genuine evaluation.
Together, these measures ensure that our final dataset of 1,095 participants reflects thoughtful and engaged responses; approximately 25% of initially recruited participants were removed after failing attention checks, minimum-time thresholds, or post-hoc quality-control criteria.
Demographic Breakdown of Study Participants
This section presents the demographic characteristics of the 1,095 participants who complete our study. To ensure ethical data collection and respect participant privacy, we provide an “I wish not to declare” option for all demographic questions, allowing participants to opt out without being forced to provide false information. The sample includes diverse representation across age groups, gender identities, countries of residence, ethnic backgrounds, and political affiliations, providing a comprehensive foundation for analyzing responses to the survey questions.
Table 3: Demographic Breakdown of Study Participants
Category Subcategory Count (%)
Total Participants
Total 1095 (100.00%)
Age
18–30 485 (44.29%)
31–40 319 (29.13%)
41–50 148 (13.52%)
51–60 94 (8.58%)
60+ 46 (4.20%)
Wish not to declare 3 (0.27%)
Gender
He/Him/His 510 (46.58%)
She/Her/Hers 549 (50.14%)
They/Them/Theirs 18 (1.64%)
Wish not to declare 18 (1.64%)
Country of Residence
Germany 525 (47.95%)
USA 563 (51.42%)
Wish not to declare 7 (0.64%)
Ethnicity
White 748 (68.31%)
Black/African American 138 (12.60%)
Asian 69 (6.30%)
Mixed 51 (4.66%)
Hispanic/Latino 28 (2.56%)
Middle Eastern/N. African 23 (2.10%)
American Indian/Alaska Native 6 (0.55%)
Wish not to declare 32 (2.92%)
Political Spectrum
Rather Liberal 343 (31.32%)
Centre 300 (27.40%)
Liberal 244 (22.28%)
Rather Conservative 110 (10.05%)
Conservative 98 (8.95%)
Political Party
Republicans 193 (17.63%)
Democrats 281 (25.66%)
AfD 28 (2.56%)
Andere 79 (7.21%)
CDU/CSU 56 (5.11%)
FDP 28 (2.56%)
Grüne 135 (12.33%)
Linke 65 (5.94%)
Piraten 11 (1.00%)
SPD 66 (6.03%)
Tier 15 (1.37%)
Wish not to declare 138 (12.60%)
Distribution of Ratings Across Value Dimensions
To provide an overview of how human raters evaluated the language model responses, we visualized the distribution of ratings across the five value dimensions(Figure 7): stereotypical bias, toxicity, emotional awareness, sensitivity, and helpfulness. Each dimension was evaluated on a 5-point Likert scale ranging from strongly disagree to strongly agree.
Figure 7: Distribution of participant ratings across five value dimensions. Ratings are on a 5-point Likert scale: strongly disagree, disagree, neutral, agree, and strongly agree.
A.5 Limitations and Future Work
Our dataset includes participants from the US and Germany, limiting cultural scope; we also recognize that both countries represent predominantly WEIRD (Western, Educated, Industrialized, Rich, and Democratic) cultures, and suggest that future work be conducted across a more diverse set of countries and cultures. We standardized political ideology using a five-point scale and harmonized ethnicity categories across countries, while acknowledging inherent cross-national differences. To ensure sufficient sample sizes, we collapsed adjacent ideological groups (e.g., “conservative” and “rather conservative”). (See Supplementary Table 3 for participant demographics) Still, conservatives, gender minorities, and older adults remained underrepresented, reflecting common online recruitment challenges. Future work should expand geographic and demographic diversity.
While using GPT-4o-mini for evaluation may miss cultural nuance, we validated its use through two human experts who independently evaluated 50 responses, achieving 85% agreement across toxicity and emotional awareness dimensions. Furthermore, our optimization analysis focused on DPO versus GRPO, leaving Proximal Policy Optimization (PPO) and Constitutional AI for future work.
Scale granularity analysis involved converting 5-point ratings to 3-point and binary formats. Although systematically applied, this process may not fully preserve semantic meaning. Future studies should directly collect ratings at different granularities and examine optimal scale designs for capturing value diversity across demographic groups.
A.6 Detailed Regression Results
This appendix presents the complete ordinal regression re- sults for all five rating dimensions analyzed in our study. For regression analysis, we collapsed the 5-point Likert scale into a 3-point scale (Agree, Neutral, Disagree) to improve model stability by increasing observations per category and to enhance interpretability of the estimated effects. Each table reports coefficient estimates, odds ratios, and significance levels for demographic predictors. Significance is denoted as: p<0.05*p<0.05, p<0.01**p<0.01, p<0.001***p<0.001.
Table 4: CLMM coefficients for Toxicity. Coefficient estimates and significance levels.
Variable Est. OR p-value Sig
Country: Germany -0.073 0.93 0.325
Male (he/him) -0.197 0.821 0.002 *
Gender: Non-binary -0.099 0.906 0.675
Age: 31-40 -0.071 0.931 0.34
Age: 41-50 -0.049 0.952 0.622
Age: 51-60 -0.095 0.909 0.422
Age: 60+ -0.072 0.93 0.657
Pol: Centre 0.098 1.103 0.183
Pol: Conservative -0.06 0.941 0.497
Ethnicity:Black/African American 0.015 1.015 0.884
Ethnicity: Hispanic/Latino 0.231 1.26 0.211
Ethnicity: Asian -0.01 0.99 0.941
Ethnicity: MENA 0.185 1.203 0.383
Ethnicity: Mixed -0.146 0.864 0.323
Table 5: CLMM coefficients for Helpfulness. Coefficient estimates and significance levels.
Variable Est. OR p-value Sig
Country: Germany -0.149 0.861 0.073
Male (he/him) 0.069 1.071 0.347
Gender: Non-binary -0.228 0.796 0.4
Age: 31-40 0.07 1.072 0.409
Age: 41-50 -0.059 0.943 0.603
Age: 51-60 -0.521 0.594 0 *
Age: 60+ -0.095 0.91 0.607
Pol: Centre -0.049 0.952 0.556
Pol: Conservative 0.004 1.004 0.967
Ethnicity:Black/African American 0.106 1.112 0.371
Ethnicity: Hispanic/Latino -0.239 0.788 0.252
Ethnicity: Asian 0.052 1.054 0.73
Ethnicity: MENA -0.387 0.679 0.11
Ethnicity: Mixed 0.012 1.013 0.94
Table 6: CLMM coefficients for Sensitivity. Coefficient estimates and significance levels.
Variable Est. OR p-value Sig
Country: Germany -0.022 0.978 0.766
Male (he/him) 0.069 1.072 0.29
Gender: Non-binary -0.274 0.76 0.257
Age: 31-40 0.044 1.045 0.558
Age: 41-50 0.01 1.01 0.921
Age: 51-60 -0.071 0.931 0.55
Age: 60+ -0.009 0.991 0.956
Pol: Centre 0.122 1.129 0.101
Pol: Conservative 0.246 1.279 0.006 *
Ethnicity:Black/African American 0.459 1.582 0 *
Ethnicity: Hispanic/Latino 0.074 1.077 0.689
Ethnicity: Asian 0.166 1.18 0.213
Ethnicity: MENA 0.133 1.143 0.527
Ethnicity: Mixed 0.063 1.066 0.671
Table 7: CLMM coefficients for Stereotypical Bias. Coefficient estimates and significance levels.
Variable Est. OR p-value Sig
Country: Germany -0.115 0.892 0.119
Male (he/him) -0.234 0.791 0 *
Gender: Non-binary -0.037 0.964 0.878
Age: 31-40 0.051 1.052 0.496
Age: 41-50 0.137 1.146 0.169
Age: 51-60 0.028 1.028 0.814
Age: 60+ -0.056 0.946 0.733
Pol: Centre 0.122 1.13 0.096
Pol: Conservative 0.038 1.039 0.67
Ethnicity:Black/African American -0.003 0.997 0.975
Ethnicity: Hispanic/Latino 0.025 1.025 0.893
Ethnicity: Asian -0.115 0.892 0.379
Ethnicity: MENA 0.06 1.062 0.778
Ethnicity: Mixed -0.017 0.983 0.909
Table 8: CLMM coefficients for Emotional Awareness. Coefficient estimates and significance levels.
Variable Est. OR p-value Sig
Country: Germany 0.075 1.078 0.333
Male (he/him) 0.012 1.012 0.858
Gender: Non-binary -0.304 0.738 0.237
Age: 31-40 0.035 1.035 0.659
Age: 41-50 -0.08 0.923 0.445
Age: 51-60 -0.103 0.902 0.407
Age: 60+ -0.002 0.998 0.99
Pol: Centre 0.065 1.067 0.4
Pol: Conservative 0.24 1.271 0.01 *
Ethnicity:Black/African American 0.362 1.436 0.001 *
Ethnicity: Hispanic/Latino 0.249 1.283 0.202
Ethnicity: Asian 0.04 1.041 0.774
Ethnicity: MENA 0.362 1.436 0.1
Ethnicity: Mixed 0.004 1.004 0.979
A.7 Alignment Assumptions and Our Framing
Our work contrasts the standard “one-model-for-all” alignment pipeline with a pluralistic alternative. Traditional alignment assumes that a small set of annotators and a fixed set of values can speak for everyone, producing a single model trained on uniform feedback. Our approach instead varies who provides the feedback and how the feedback is encoded. By using diverse raters and testing different rating scales, aggregation methods and optimisation choices, we show that alignment outcomes change depending on these social and technical decisions. This demonstrates that alignment is not inherently universal and that accounting for variation leads to models that better reflect diverse perspectives.
Figure 8: Traditional monolithic alignment assumes universal values of helpfulness, honesty, and harmlessness (left), whereas our approach varies rater demographics and technical choices (rating scales, disagreement aggregation, and optimization methods) to reflect diverse perspectives (right).
Appendix B Supplementary B
B.1 Experimental Design Overview
Figure 10 provides a systematic overview of our four fine-tuning experiments, showing how we systematically vary key parameters while controlling others. Each experiment isolates specific factors affecting pluralistic alignment: demographic stratification (Experiment 1), rating scale granularity (Experiment 2), disagreement handling methods (Experiment 3), and optimization algorithms (Experiment 4). This design allows us to assess the individual and comparative effects of demographic composition, technical choices, and methodological decisions on model alignment performance.
Figure 9: Evaluation prompt specifications. Standardized protocols for toxicity assessment (continuous 0.0-1.0 scale) and emotional awareness evaluation (binary 0-1 scale) using GPT-4o-mini as evaluator.
B.2 Complete Model and Experimental Specifications
All models used in this study were available on Hugging Face throughout the fine-tuning period (March–September 2025), ensuring stable and reproducible training. The selection captures variation in size, objectives, and safety properties to support comparative analysis across architectures. Table 9 lists the models included.
Table 9: Model Architecture Overview. Primary specifications for all evaluation models.
Model Params Specialization
Wizard-Vicuna-7B (cognitivecomputations) 7B Instruction + Uncensored
Llama-3-8B-Lexi (Orenguteng) 8B General-purpose + Uncensored
Deep-Reasoning-Llama-3.2 (DavidAU) 3B Reasoning + Uncensored
Qwen3-14B Abliterated (mlabonne) 14B General-purpose + Abliterated
Guanaco-3B-Uncensored-v2 (Fredithefish) 3B Instruction-tuned + Uncensored
Llama3.2-1B-Uncensored (xdrshjr) 1B General-purpose + Uncensored
WizardVicuna-3B (heegyu) 3B Instruction + Uncensored
B.3 Evaluation Methodology
We evaluate alignment performance using OpenAI’s GPT-4o-mini API to score model responses on Toxicity (continuous 0.0-1.0 scale) and Emotional Awareness dimensions (binary 0-1). We use two different scales because toxicity, following the continuous scoring framework of (Jigsaw and Google ), varies by degree, whereas emotional awareness, consistent with Sharma et al. (2020), is expressed in threshold-like categories and is therefore more reliably captured as a binary judgement. To validate this automated approach, we conduct a reliability study where two independent human experts assess 50 randomly selected responses using identical criteria. We achieve an average agreement rate of 0.85 between human evaluators and automated scores across both dimensions, confirming the reliability of our evaluation approach. Our automated evaluation employs GPT-4o-mini with temperature=0 for consistent scoring. Figure 9 shows the prompts used for each alignment dimension.
B.4 Alignment Optimization Methods
This subsection summarizes the optimization procedures used in our experiments and provides the mathematical formulations for the preference learning objectives applied in both DPO and GRPO.
Single-Objective DPO
Classical Direct Preference Optimization (DPO) on a single dimension (e.g., toxicity) minimizes the pairwise logistic loss:
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}\left[\log\sigma\left(s_{\theta}(y^{+};x)-s_{\theta}(y^{-};x)\right)\right],$ (2)
where
$\sigma(z)$ $=\frac{1}{1+e^{-z}},$ (3)
$s_{\theta}(y;x)$ $=\log\pi_{\theta}(y\mid x)-\log\pi_{\mathrm{ref}}(y\mid x).$ (4)
Group Relative Policy Optimization (GRPO)
For each prompt $q$ and model-generated output $o_{i}$, the reward function $R(q,o_{i})$ assigns a scalar value $r_{i}\in[0,1]$ based on a Likert-style human rating. We use two distinct mappings depending on whether the evaluation dimension is negatively or positively oriented. We implement GRPO using the Hugging Face TRL library’s GRPOTrainer (Zhihong Shao 2024).
Negative-oriented dimensions. For $d\in\{\text{toxicity},\ \text{stereotypical bias}\}$, agreement indicates undesirable behavior. The reward function is:
[…]
Positive-oriented dimensions. For $d\in\{\text{sensitivity},\ \text{helpfulness},\ \text{emotional awareness}\}$, agreement indicates desirable behavior. The reward function is:
[…]
Computing the loss. The GRPO objective maximizes the advantage while keeping the policy close to the reference model. The loss is:
[Uncaptioned image] (7)
The scalar reward is then used within GRPO by first computing the group-normalized reward
$\tilde{r}_{i}=\frac{r_{i}-\mathrm{mean}(r_{1},\dots,r_{G})}{\mathrm{std}(r_{1},\dots,r_{G})},$ (8)
which defines the per-token advantage
$\hat{A}_{i,t}=\tilde{r}_{i},\qquad\forall t\in\{1,\dots,T_{i}\}.$ (9)
Multi-Objective DPO
To optimize two objectives simultaneously (e.g., toxicity and emotional awareness), we pool all accept/reject pairs-”Accept” refers to the response rated higher by participants on the target dimension, and ”reject” is the lower-rated response- from both dimensions into one dataset and apply the same loss:
$\mathcal{L}_{\text{multi-DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}\big[\log\sigma\big(s_{\theta}(y^{+}\mid x)\\ -s_{\theta}(y^{-}\mid x)\big)\big]$ (10)
where each $(y^{+},y^{-})$ may originate from either objective’s judgments.
DPO with Different Aggregation Strategies
We use the same DPO loss but vary how we form $(y^{+},y^{-})$ from the raw Likert annotations:
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\mathbb{E}_{(x,y^{+},y^{-})}\left[\log\sigma\left(s_{\theta}(y^{+})-s_{\theta}(y^{-})\right)\right].$ (11)
We construct five versions of the dataset using different aggregation strategies to examine how disagreement resolution affects alignment outcomes:
  • All Ratings: Includes all individual rater responses without aggregation. Each accept/reject judgment is used as a training pair.
  • Majority Vote: Assigns the most frequent label across raters for each prompt-response pair. Ties are broken randomly.
  • Full Consensus: Retains only examples where all raters unanimously agreed on the label. This method emphasizes certainty but results in fewer examples.
  • Random Selection: Randomly samples one rater’s label per item. This simulates a lightweight annotation strategy with minimal aggregation.
  • Create Average: Converts Likert ratings to numeric scores, computes the mean per item, and rounds to the nearest valid category.
Each strategy encodes a different assumption about how best to handle disagreement, enabling us to assess the impact of aggregation on preference optimization.
B.5 Random-Effects Meta-analysis Formulation
We employ the DerSimonian–Laird random-effects estimator (DerSimonian and Laird 1986) to synthesize results across model architectures. For each condition, we pool effect sizes $\theta_{i}$ (differences between fine-tuned and control models) using random-effects weights:
$\hat{\theta}=\frac{\sum_{i=1}^{k}w_{i}^{*}\theta_{i}}{\sum_{i=1}^{k}w_{i}^{*}},\quad w_{i}^{*}=\frac{1}{\sigma_{i}^{2}+\hat{\tau}^{2}}$ (12)
where $\sigma_{i}^{2}$ is the within-model variance and $\hat{\tau}^{2}$ is the estimated between-model heterogeneity.
For pairwise comparisons between conditions $(j,k)$, we compute differences $\hat{\delta}_{jk}=\hat{\theta}_{j}-\hat{\theta}_{k}$ with associated variance:
$\text{Var}(\hat{\delta}_{jk})=\frac{1}{\sum w_{i}^{*(j)}}+\frac{1}{\sum w_{i}^{*(k)}}$ (13)
Statistical significance is determined based on 95% confidence intervals that exclude zero.
B.6 MMLU Evaluation Results
We evaluated seven models across 17 fine-tuning runs using 1000 MMLU questions. Baseline accuracy ranged from 23.6% (Guanaco-3B) to 56.9% (Llama-3-8B, Qwen3-14B). Qwen3-14B was the most sensitive to fine-tuning, with White EA DPO improving accuracy by +4.5% (61.4%) and consensus selection reducing it by -5.5%. Rating-scale effects varied by model: binary scales aided Wizard-Vicuna-7B (+1.0%) but harmed Deep-Reasoning (-1.9%). Data-selection strategies were likewise model-dependent, with random sampling best for Guanaco-3B (+1.2%) and majority-vote selection optimal for Llama models. Conservative-aligned training had mixed influence (+1.5% for Wizard-Vicuna-7B, negligible elsewhere).
The MMLU results are not an evaluation of our fine-tuning task. MMLU covers general-knowledge and reasoning skills (math, science, history), which are unrelated to toxicity or emotional-awareness training. We report MMLU only to show baseline model capability. Some models, such as Wizard-Vicuna-7B and Guanaco-3B, have low general-knowledge scores, so their limitations come from the base model rather than from our fine-tuning. The higher-performing models (for example, Qwen3-14B and Llama-3-8B) are therefore easier to interpret because they start from a stronger overall capability.
Table 10: MMLU Accuracy:Deep-Reasoning-Llama-3.2-Instruct-uncensored-3B
Variant Accuracy (%) vs Base (%)
Base Model 52.9 baseline
Demographics
Male Toxic DPO 51.0 -1.9
Female Toxic DPO 50.9 -2.0
Black EA DPO 52.3 -0.6
White EA DPO 52.1 -0.8
Conservative EA DPO 52.8 -0.1
Liberal EA DPO 53.1 +0.2
Rating Scales
5-points Toxic DPO 52.3 -0.6
3-points Toxic DPO 52.3 -0.6
Binary Toxic DPO 51.0 -1.9
Data Selection
Full Dataset (5-points) 52.3 -0.6
Majority Vote DPO 50.5 -2.4
Create Average DPO 51.7 -1.2
Keep Random DPO 51.1 -1.8
Full Consensus DPO 52.7 -0.2
Other Variants
EA DPO 52.5 -0.4
Toxic EA DPO 51.9 -1.0
Toxic EA GRPO 52.9 +0.0
Table 11: MMLU Accuracy: Qwen3-14B-abliterated
Variant Accuracy (%) vs Base (%)
Base Model 56.9 baseline
Demographics
Male Toxic DPO 57.7 +0.8
Female Toxic DPO 59.6 +2.7
Black EA DPO 58.8 +1.9
White EA DPO 61.4 +4.5
Conservative EA DPO 57.5 +0.6
Liberal EA DPO 60.3 +3.4
Rating Scales
5-points Toxic DPO 59.4 +2.5
3-points Toxic DPO 58.9 +2.0
Binary Toxic DPO 56.3 -0.6
Data Selection
Full Dataset (5-points) 59.4 +2.5
Majority Vote DPO 53.3 -3.6
Create Average DPO 56.7 -0.2
Keep Random DPO 54.3 -2.6
Full Consensus DPO 51.4 -5.5
Other Variants
EA DPO 60.1 +3.2
Toxic EA DPO 59.1 +2.2
Toxic EA GRPO 57.1 +0.2
Table 12: MMLU Accuracy: Wizard-Vicuna-7B-Uncensored
Variant Accuracy (%) vs Base (%)
Base Model 31.5 baseline
Demographics
Male Toxic DPO 30.7 -0.8
Female Toxic DPO 31.4 -0.1
Black EA DPO 31.8 +0.3
White EA DPO 30.5 -1.0
Conservative EA DPO 33.0 +1.5
Liberal EA DPO 30.3 -1.2
Rating Scales
5-points Toxic DPO 29.7 -1.8
3-points Toxic DPO 31.5 +0.0
Binary Toxic DPO 32.5 +1.0
Data Selection
Full Dataset (5-points) 29.7 -1.8
Majority Vote Toxic DPO 31.0 -0.5
Create Average Toxic DPO 31.4 -0.1
Keep Random Toxic DPO 31.5 +0.0
Full Consensus Toxic DPO 31.5 +0.0
Other Variants
EA DPO 29.2 -2.3
Toxic EA DPO 31.0 -0.5
Toxic EA GRPO 31.5 +0.0
Table 13: MMLU Accuracy: Llama-3-8B-Lexi-Uncensored
Variant Accuracy (%) vs Base (%)
Base Model 56.9 baseline
Demographics
Male Toxic DPO 57.2 +0.3
Female Toxic DPO 56.3 -0.6
Black EA DPO 57.4 +0.5
White EA DPO 57.5 +0.6
Conservative EA DPO 55.8 -1.1
Liberal EA DPO 54.2 -2.7
Rating Scales
5-points Toxic DPO 56.4 -0.5
3-points Toxic DPO 55.3 -1.6
Binary Toxic DPO 57.2 +0.3
Data Selection
Full Dataset (5-points) 56.4 -0.5
Majority Vote Toxic DPO 57.6 +0.7
Create Average Toxic DPO 56.8 -0.1
Keep Random Toxic DPO 57.0 +0.1
Full Consensus Toxic DPO 57.0 +0.1
Other Variants
EA DPO 56.5 -0.4
Toxic EA DPO 55.6 -1.3
Toxic EA GRPO 56.9 +0.0
Table 14: MMLU Accuracy: Guanaco-3B-Uncensored
Variant Accuracy (%) vs Base (%)
Base Model 23.6 baseline
Demographics
Male Toxic DPO 22.8 -0.8
Female Toxic DPO 24.1 +0.5
Black EA DPO 23.7 +0.1
White EA DPO 24.0 +0.4
Conservative EA DPO 22.0 -1.6
Liberal EA DPO 23.8 +0.2
Rating Scales
5-points Toxic DPO 23.7 +0.1
3-points Toxic DPO 23.1 -0.5
Binary Toxic DPO 23.8 +0.2
Data Selection
Full Dataset (5-points) 23.7 +0.1
Majority Vote Toxic DPO 22.9 -0.7
Create Average Toxic DPO 23.2 -0.4
Keep Random Toxic DPO 24.8 +1.2
Full Consensus Toxic DPO 23.2 -0.4
Other Variants
EA DPO 23.3 -0.3
Toxic EA DPO 23.6 +0.0
GRPO Fine-tuned 23.6 +0.0
Table 15: MMLU Accuracy: WizardVicuna-Uncensored-3B
Variant Accuracy (%) vs Base (%)
Base Model 24.8 baseline
Demographics
Male Toxic DPO 22.8 -2.0
Female Toxic DPO 24.5 -0.3
Black EA DPO 24.8 +0.0
White EA DPO 21.9 -2.9
Conservative EA DPO 23.2 -1.6
Liberal EA DPO 23.5 -1.3
Rating Scales
5-points Toxic DPO 26.0 +1.2
3-points Toxic DPO 23.1 -1.7
Binary Toxic DPO 22.2 -2.6
Data Selection
Full Dataset (5-points) 26.0 +1.2
Majority Vote Toxic DPO 23.3 -1.5
Create Average Toxic DPO 25.0 +0.2
Keep Random Toxic DPO 22.9 -1.9
Full Consensus Toxic DPO 21.8 -3.0
Other Variants
EA DPO 21.1 -3.7
Toxic EA DPO 25.6 +0.8
GRPO DPO 24.9 +0.1
Table 16: MMLU Accuracy: llama3.2-1b-uncensored-5000-8epoch-lora
. Variant Accuracy (%) vs Base (%) Base Model 40.5 baseline Demographics Male Toxic DPO 40.1 -0.4 Female Toxic DPO 41.0 +0.5 Black EA DPO 40.4 -0.1 White EA DPO 40.2 -0.3 Conservative EA DPO 40.6 +0.1 Liberal EA DPO 38.9 -1.6 Rating Scales 5-points Toxic DPO 40.0 -0.5 5-points Toxic DPO 41.2 +0.7 Binary Toxic DPO 40.4 -0.1 Data Selection Full Dataset (5-points) 40.0 -0.5 Majority Vote Toxic DPO 41.3 +0.8 Create Average Toxic DPO 40.4 -0.1 Keep Random Toxic DPO 41.2 +0.7 Full Consensus Toxic DPO 41.8 +1.3 Other Variants EA DPO 40.2 -0.3 Toxic EA DPO 41.0 +0.5 Toxic EA GRPO 40.7 +0.2
Figure 10: Experimental design overview. Systematic investigation of pluralistic alignment factors across four controlled experiments varying demographic stratification, scale granularity, aggregation methods, and optimization algorithms.